# DemoDiffusion: One-Shot Human Imitation using pre-trained Diffusion Policy

### v1

@@ Abstract
We propose DemoDiffusion, a simple and scalable method for enabling robots to perform manipulation tasks in natural environments by imitating a single human demonstration. Our approach is based on two key insights. First, the hand motion in a human demonstration provides a useful prior for the robot’s end-effector trajectory, which we can convert into a rough open-loop robot motion trajectory via kinematic retargeting. Second, while this retargeted motion captures the overall structure of the task, it may not align well with plausible robot actions in-context. To address this, we leverage a pre-trained generalist diffusion policy to modify the trajectory, ensuring it both follows the human motion and remains within the distribution of plausible robot actions. Our approach avoids the need for online reinforcement learning or paired human-robot data, enabling robust adaptation to new tasks and scenes with minimal manual effort. Experiments in both simulation and real-world settings show that DemoDiffusion outperforms both the base policy and the retargeted trajectory, enabling the robot to succeed even on tasks where the pre-trained generalist policy fails entirely. Project page: https://demodiffusion.github.io/
Keywords: Diffusion Policy, One-shot Imitation, Human to Robot
Figure 1: Overview of  DemoDiffusion. We show how generalist pre-trained diffusion policies can be used for following a generic human demonstration showing a manipulation task during deployment. Our real-world manipulation results encompass a wide diversity of manipulation tasks involving everyday objects.
1 Introduction
@@
How do we build robot manipulation systems that can be readily deployed in unstructured human environments? One possible answer is to learn ‘generalist’ policies that are capable of accomplishing any generic task (specified via some description e.g., language or image goal) in any environment. Indeed, there is a general optimism about this paradigm, reflected in several ongoing efforts to collect large-scale demonstration datasets to train such policies. While these efforts have led to impressive results in the form of unified policies capable of performing diverse tasks [1, 2, 3, 4, 5], these policies still struggle to perform meaningfully when deployed zero-shot to novel environments or asked to perform unseen tasks. Deployment thus often requires additional fine-tuning using task-and-scene-specific robot demonstrations [1], but this is a non-trivial overhead as collecting robot demonstrations in the real world can be time-consuming and beyond the expertise of an average user. In this work, we propose an alternate deployment mechanism – leveraging a pre-trained generalist policy to perform a task via imitating a single human demonstration.
@@
We are of course not the first to consider this goal of allowing robots to imitate any given human demonstration. One common approach [6, 7, 8, 9] is to instantiate this as kinematic retargeting task and compute open-loop robot actions that maximize a manually defined similarity between the achieved robot end-effector configuration and observed human hand poses (e.g., matching locations of fingertips). However, the human-robot embodiment mismatch makes it difficult for the retargeted actions to achieve precisely the same effects as the human ones, and the open-loop execution further makes the approach brittle to noise and scene variations. Another line of work for human imitation attempts to learn robot policies through online reinforcement learning [10, 11, 12, 13, 14], where the human demonstration helps define reward functions. While this can overcome the embodiment gap, such test-time online RL requires hours of online interaction and resets - making it difficult to adopt such methods for generic real-world manipulation tasks, especially in safety-critical scenarios. We instead seek to develop an approach for demonstration following that, akin to kinematic retargeting, can be deployed ‘one-shot’ i.e., without any test-time training while still benefitting from generic learning-based priors for more precise closed-loop interaction.
@@
Our approach builds on the insight that pre-trained diffusion policies can act as priors for robot action. Inspired by prior work in leveraging pre-trained diffusion models for image editing [15], we present DemoDiffusion–a formulation to utilize diffusion policy trained on robot interaction data for synthesizing coherent robot actions from a human demonstration. Specifically, we first perform kinematic retargeting by extracting human hand poses from the demonstration and obtain an open-loop robot action trajectory. While this trajectory is typically suboptimal due to embodiment differences and lack of closed-loop feedback, it serves as an effective initialization that can be improved via a diffusion policy. We do so by injecting gaussian noise and appling the pre-trained diffusion policy to iteratively denoise the trajectory conditioned on robot observations, yielding a refined, executable sequence of robot actions. DemoDiffusion thus enables the robot to use the pre-trained policy as a prior and adapt the human-derived trajectory to its own embodiment and environment in a closed-loop manner.
@@
In summary, we propose DemoDiffusion: a framework for robotic manipulation that allows a robot to perform generic tasks in natural environments by following a human demonstration with guidance from a pre-trained diffusion policy. Importantly, our framework doesn’t require any robot demonstration of the target task in the target environment and doesn’t require any online interaction or fine-tuning. Our experiments across simulation and real-world environments show that DemoDiffusion surpasses the performance of both, the base policy and retargeted trajectory, even allowing the robot to perform tasks where pre-trained generalist policy completely fails.
2 Related Work
Generalist Manipulation Policies.
There has been a growing trend in developing ‘generalist’ robotic policies that can perform multiple tasks based on a specified goal in the form of an image or a language instruction. Collecting large-scale robot interaction datasets [16, 17] combined with behavior cloning and vision-language pre-training [5, 2, 18, 1, 4, 3] is the predominant recipe for training such multi-task policies. However, since collecting large-scale robot interaction datasets in natural settings like homes and offices is challenging due to operational constraints, current best models still struggle to perform manipulation tasks zero-shot and thus are not yet deployable in-the-wild. Our work enables using such generalist diffusion policies for following a human demonstration, thereby being able to perform tasks that the pre-trained policy might be unable to zero-shot.
Robotic Manipulation with Non-Robot Datasets.
Instead of requiring large-scale robot interaction datasets, a growing body of works have begun utilizing human videos and large-scale web videos for robotics. A lot of these approaches have been enabled by recent advances in computer vision for representation learning [19, 20, 21], predicting tracks [22, 23], and reconstructing hand-object interactions from monocular videos [24, 25, 26, 27, 27]. Earlier works learned self-supervised visual representations from such non-robotic datasets [28, 29, 30] that can serve as the visual backbone of robotic policies [31, 32, 33, 34, 35, 36]. Recent works in this paradigm predict manipulation-relevant cues from web videos in the form of motion trajectories [37, 38] and object affordances [39, 40] and combine these predictive models with a limited amount of robot interaction data for training policies via conditional behavior cloning. Our work is orthogonal to these approaches in that instead of using human videos for training, we require a single human demonstration during deployment for guiding a pre-trained generalist policy to perform a new task.
One-Shot Imitation from Robot Demonstration.
One-shot imitation aims to enable robots to perform a new task with guidance from a single demonstration. Prior works have explored this in the context of a robot demonstration provided at test time, via training demonstrateion-conditioned policies [41]. Recent works have developed learning-based visual servoing systems [42] and algorithms for identifying object invariances [43] to replicate the robot end-effector actions from the demonstration onto novel configurations of objects. However, requiring a robot demonstration for imitation is restrictive for ubiquitous deployment as end-users might find it challenging to tele-operate the robot to collect demonstrations in natural human environments.
One-Shot Imitation from Human Demonstration.
To imitate a human demonstration for robot manipulation, a straightforward approach is kinematic retargeting of the human hand pose to the robot end-effector pose per time-step, as demonstrated by some recent works [6, 7]. Although simple to implement, the human-robot embodiment mismatch typically introduces errors in the retargeting and the open-loop execution is brittle to object pose variations during deployment. To remedy this, another line of works seek to learn this retargeting implicitly by training a closed-loop policy from paired human-robot datasets [44, 45] enabling more flexible deployments. Since such paired datasets are difficult to collect, other works [10, 11, 46, 12] perform reinforcement learning with reward functions derived from video comparison. However, such methods require online interaction and resets, limiting their practicality. We propose an alternate paradigm: given the human demonstration, we refine retargeted human behavior using a pre-trained robot policy, avoiding brittle replays and enabling closed-loop control. This formulation doesn’t rely on expensive paired human-robot data collection and also doesn’t require cumbersome test-time fine-tuning via interaction.
3 Method
@@
Figure 2: Retargeted human hand trajectory to closed-loop robot action sequence, for the task $\mathcal{T}$: “close the laptop”. The dotted line shows the trajectory of robot end-effector poses after kinematic retargeting. The olive contour plot depicts the distribution of trajectories from a pre-trained diffusion policy. Given a kinematic retargeting, we first perturb it with Gaussian noise and progressively remove the noise by simulating the reverse SDE with the diffusion policy. This process gradually projects a potentially unfeasible but approximately correct retargeting to the manifold of plausible robot actions that can perform real-world manipulation, in this case closing the laptop without missing the edge.
@@
We target the problem of one-shot visual imitation. Given a human demonstration $\mathbf{D}$ depicting a manipulation task with description $\mathcal{T}$, we want a robot manipulator to perform the same task. Unlike prior work that relies on test-time RL training [10] or paired human-robot datasets [44], we wish to enable such imitation ‘one-shot’ assuming access to a closed-loop diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ pre-trained on some broad robot interaction dataset $\mathcal{D}_{robot}$. Our assumption on the human demonstration $\mathbf{D}$ is also very minimal: it can either be an RGBD video or a multi-view video of a human doing a task, such that 3D hand poses per-timestep can be reliably extracted from it.
3.1 Overview
@@
Our approach is based on two key insights. The first is that the trajectory of the hand pose in the human demonstration $\mathbf{D}$ provides useful information of the approximate trajectory the robot end-effector should follow, and we can perform kinematic retargeting of the hand trajectory $\{\mathbf{h}_{t}\}^{T}_{t=0}$ to an open-loop robot end-effector trajectory $\{\hat{\mathbf{a}}_{t}\}^{T}_{t=0}$. The second insight is that the kinematically retargeted robot trajectory has the correct form of motion, but these actions may not be very precise in the distribution of plausible robot actions given the current observation. A diffusion policy models this likelihood, and we can use a pre-trained ‘generalist’ diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ to refine the retargeted robot actions in a closed loop manner, thus inferring actions that, while still similar to the human demonstration, are more likely under the policy and better aligned with the robot embodiment (see Fig. 2). We describe both these steps in detail below.
@@
3.2 Kinematic retargeting of Human Hand Trajectories to Robot End-Effector Poses
@@
Figure 3: Dexterous Grasping Results in simulation. On the left, we show a human demonstration followed by a rollout with DemoDiffusion for dexterous grasping. On the right, we plot the average and standard deviation of success rates over 3 seeds for dexterous grasping, corresponding to different levels of the diffusion step $s^{*}$. Here, $s^{*}/S=0$ corresponds to kinematic retargeting and $s^{*}/S=1$ corresponds to the robot policy.
3.3 Closed-Loop Denoising of Robot Actions with a Pre-Trained Diffusion Policy
A diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ [48, 49] pre-trained on diverse offline robot interaction data reliably models the distribution of plausible robot actions $\mathbf{a}_{t}$ given previous observations $\mathbf{o}_{\leq t}$. To predict actions $\mathbf{a}_{t}$, diffusion policies start with Gaussian noise $\tilde{\mathbf{a}}_{t}^{(S)}\sim\mathcal{N}(0,I)$ and iteratively denoise it at different diffusion steps $s$:
@@ 3.3 Closed-Loop Denoising of Robot Actions with a Pre-Trained Diffusion Policy
$\tilde{\mathbf{a}}_{t}^{(s^{*})}=\sqrt{\alpha_{s^{*}}}\hat{\mathbf{a}}_{t}+\sqrt{1-\alpha_{s^{*}}}\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ (2)
Here $\alpha$ corresponds to the diffusion schedule of the pre-trained policy. This procedure, inspired by SDEdit [15] which adopted a similar approach for image editing, relies on the assumption that $\{\hat{\mathbf{a}}_{t}\}_{t=0}^{T}$ is an approximate version of the ideal action trajectory that should be executed by the robot, and it potentially lies outside the distribution of feasible actions under the pre-trained diffusion policy $\bar{\pi}_{\theta}$. The diffusion policy $\bar{\pi}_{\theta}$ then performs iterative denoising steps, conditioned on the robot’s observations $o_{\leq t}$, to refine this noisy trajectory into feasible robot actions, based on equation (2).
After $s^{*}$ denoising steps, the final output $\mathbf{a}_{t}=\tilde{\mathbf{a}}_{t}^{(0)}$ is deployed on the robot. Importantly, this process is carried out in a closed-loop manner: the policy uses real-time observations from the cameras in the scene to iteratively improve its predictions, thereby compensating for embodiment mismatch and external perturbations (e.g., object slippage or occlusion). The key hyperparameter in this process for DemoDiffusion is the diffusion step $s^{*}$ that trades off between the faithfulness to the demonstration and the likelihood under the robot policy – in the limit $s^{*}=S$ we recover a rollout from the base policy $\bar{\pi}_{\theta}$ and in the limit $s^{*}=0$ we recover the kinematically retargeted trajectory $\{\hat{\mathbf{a}}_{t}\}_{t=0}^{T}$.
Table 1: Simulated Dexterous Grasping Results. The numbers show the average of success rates over 3 seeds. We use $s^{*}/S=0.2$, where the robot policy uses $S=1000$DemoDiffusion consistently outperforms both baselines, across all groups.
Method Small Medium Large Average
Robot Policy 25.4 26.6 27.4 26.5
Kinematic Retargeting 2.6 0.7 1.5 1.6
DemoDiffusion 31.8 31.6 29.6 31.0
4 Experiments
We evaluate our approach on dexterous grasping in simulation and across diverse real-world tasks comprising of prehensile and non-prehensile table-top manipulation. Through experiments, we aim to understand the following research questions:

### v2

@@ Abstract
We propose DemoDiffusion, a simple method for enabling robots to perform manipulation tasks by imitating a single human demonstration, without requiring task-specific training or paired human-robot data. Our approach is based on two insights. First, the hand motion in a human demonstration provides a useful prior for the robot’s end-effector trajectory, which we can convert into a rough open-loop robot motion trajectory via kinematic retargeting. Second, while this retargeted motion captures the overall structure of the task, it may not align well with plausible robot actions in-context. To address this, we leverage a pre-trained generalist diffusion policy to modify the trajectory, ensuring it both follows the human motion and remains within the distribution of plausible robot actions. Unlike approaches based on online reinforcement learning or paired human-robot data, our method enables robust adaptation to new tasks and scenes with minimal effort. In real-world experiments across 8 diverse manipulation tasks, DemoDiffusion achieves 83.8% average success rate, compared to 13.8% for the pre-trained policy and 52.5% for kinematic retargeting, succeeding even on tasks where the pre-trained generalist policy fails entirely. Project page: https://demodiffusion.github.io/
I Introduction
@@
How do we build robot manipulation systems that can be readily deployed in unstructured human environments? One possible answer is to learn ‘generalist’ policies that are capable of accomplishing any generic task (specified via some description e.g. language or image goal) in any environment. Indeed, there is a general optimism about this paradigm, reflected in several ongoing efforts to collect large-scale demonstration datasets to train such policies. While these efforts have produced unified policies capable of diverse tasks [5, 62, 23, 27, 6], these policies still struggle to perform meaningfully when deployed zero-shot to novel environments or asked to perform unseen tasks. Deployment thus often requires additional fine-tuning using task-and-scene-specific robot demonstrations [5], but this is a non-trivial overhead as collecting robot demonstrations in the real world can be time-consuming and beyond the expertise of an average user. In this work, we propose an alternate deployment mechanism – leveraging a pre-trained generalist policy to perform a task via imitating a single human demonstration.
@@
We are of course not the first to consider this goal of allowing robots to imitate any given human demonstration. One common approach [29, 40, 54, 8] is to instantiate this as a kinematic retargeting task and compute open-loop robot actions that maximize a manually defined similarity between the achieved robot end-effector configuration and observed human hand poses (e.g. matching locations of fingertips). However, the human-robot embodiment mismatch makes it difficult for the retargeted actions to achieve precisely the same effects as the human ones, and the open-loop execution further makes the approach brittle to noise and scene variations. Another line of work for human imitation attempts to learn robot policies through online reinforcement learning [1, 19, 9, 31, 32], where the human demonstration helps define reward functions. While this can overcome the embodiment gap, such test-time online RL requires hours of online interaction and resets, making it difficult to adopt for generic real-world manipulation tasks, especially in safety-critical scenarios. We seek an approach that, like kinematic retargeting, can be deployed one-shot (i.e. without test-time training), but still benefits from learned priors for precise closed-loop interaction.
@@
Our approach builds on the insight that pre-trained diffusion policies can act as priors for robot action. Inspired by prior work in leveraging pre-trained diffusion models for image editing [37], we present DemoDiffusion – a formulation to utilize a diffusion policy trained on robot interaction data for synthesizing coherent robot actions from a human demonstration. Specifically, we first perform kinematic retargeting by extracting human hand poses from the demonstration and obtain an open-loop robot action trajectory. While this trajectory is typically suboptimal due to embodiment differences and lack of closed-loop feedback, it serves as an effective initialization that can be improved via a diffusion policy. We do so by injecting Gaussian noise and applying the pre-trained diffusion policy to iteratively denoise the trajectory conditioned on robot observations, yielding a refined, executable sequence of robot actions. DemoDiffusion thus enables the robot to use the pre-trained policy as a prior and adapt the human-derived trajectory to its own embodiment and environment in a closed-loop manner.
@@
In summary, we propose DemoDiffusion: a framework for robotic manipulation that allows a robot to perform generic tasks in natural environments by following a human demonstration with guidance from a pre-trained diffusion policy. Importantly, our framework doesn’t require any robot demonstration of the target task within the target environment, online interaction, or fine-tuning. Our experiments across simulation and real-world environments show that DemoDiffusion surpasses both the base policy and retargeted trajectory, achieving 83.8% success across 8 real-world tasks – even those where the pre-trained generalist policy fails entirely.
II Related Work
@@
Figure 2: Retargeted human hand trajectory to closed-loop robot action sequence, for the task $\mathcal{T}$: “shut down the laptop”. The dotted line shows the trajectory of robot end-effector poses after kinematic retargeting. The olive contour plot depicts the distribution of trajectories from a pre-trained diffusion policy. Given a kinematic retargeting, we first perturb it with Gaussian noise and progressively remove the noise by simulating the reverse SDE with the diffusion policy. This process gradually projects a potentially unfeasible but approximately correct retargeting to the manifold of plausible robot actions that can perform real-world manipulation, in this case closing the laptop without missing the edge.
Generalist Manipulation Policies. There has been a growing trend in developing ‘generalist’ robotic policies that can perform multiple tasks based on a specified goal in the form of an image or a language instruction. Collecting large-scale robot interaction datasets [16, 53] combined with behavior cloning and vision-language pre-training [6, 62, 4, 5, 27, 23, 21] is the predominant recipe for training such multi-task policies. However, since collecting large-scale robot interaction datasets in natural settings like homes and offices is challenging due to operational constraints, current best models still struggle to perform manipulation tasks zero-shot and thus are not yet deployable in-the-wild. Our work enables using such generalist diffusion policies for following a human demonstration, enabling it to perform tasks that the pre-trained policy struggles to zero-shot.
Robotic Manipulation with Non-Robot Datasets. Instead of requiring large-scale robot interaction datasets, a growing body of works have begun utilizing human videos and large-scale web videos for robotics. These approaches have been enabled by recent advances in computer vision for representation learning [20, 15, 61], predicting tracks [24, 14], and reconstructing hand-object interactions from monocular videos [47, 44, 56, 57, 57]. Earlier works learned self-supervised visual representations from such non-robotic datasets [11, 17, 18], which can serve as the visual backbone of robotic policies [39, 41, 25, 55, 35, 34]. Recent works in this paradigm predict manipulation-relevant cues from web videos in the form of motion trajectories [3, 48, 51] and object affordances [50, 2, 36, 49] and combine these predictive models with a limited amount of robot interaction data for training policies via conditional behavior cloning. Our work is orthogonal to these approaches in that instead of using human videos for training, we require a single human demonstration during deployment for guiding a pre-trained generalist policy to perform a new task.
One-Shot Imitation from Robot Demonstration. One-shot imitation aims to enable robots to perform a new task with guidance from a single demonstration. Prior works have explored this in the context of a robot demonstration provided at test time, via training demonstration-conditioned policies [12]. Recent works have developed learning-based visual servoing systems [52] and algorithms for identifying object invariances [60] to replicate the robot end-effector actions from the demonstration onto novel configurations of objects. However, requiring a robot demonstration for imitation is restrictive for ubiquitous deployment as end-users might find it challenging to tele-operate the robot to collect demonstrations in natural human environments.
One-Shot Imitation from Human Demonstration. To imitate a human demonstration for robot manipulation, a straightforward approach is kinematic retargeting of the human hand pose to the robot end-effector pose per time-step, as demonstrated by some recent works [29, 40, 28]. Although simple to implement, the human-robot embodiment mismatch typically introduces errors in the retargeting and the open-loop execution is brittle to object pose variations during deployment. To remedy this, another line of works seek to learn this retargeting implicitly by training a closed-loop policy from paired human-robot datasets [22, 42] enabling more flexible deployments. Since such paired datasets are difficult to collect, other works [1, 19, 33, 9] perform reinforcement learning with reward functions derived from video comparison. However, such methods require online interaction and resets, limiting their practicality. We propose an alternate paradigm: given the human demonstration, we refine retargeted human behavior using a pre-trained robot policy, avoiding brittle replays and enabling closed-loop control. This formulation avoids both expensive paired human-robot data collection and cumbersome test-time fine-tuning via interaction.
III Method
@@
We target the problem of one-shot visual imitation. Given a human demonstration $\mathbf{D}$ depicting a manipulation task with description $\mathscr{T}$, we want a robot manipulator to perform the same task. Unlike prior work that relies on test-time RL training [1] or paired human-robot datasets [22], we wish to enable such imitation ‘one-shot’ assuming access to a closed-loop diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ pre-trained on some broad robot interaction dataset $\mathcal{D}_{robot}$. Our assumption on the human demonstration $\mathbf{D}$ is also very minimal: it can either be an RGBD video or a multi-view video of a human doing a task, such that 3D hand poses per-timestep can be reliably extracted from it.
Figure 3: Dexterous Grasping Results in Simulation. (Left) A human demonstration and DemoDiffusion rollout for dexterous grasping. (Right) Success rates (mean $\pm$ std over 3 seeds) as a function of diffusion step $s^{*}$. Here, $s^{*}/S=0$ corresponds to kinematic retargeting and $s^{*}/S=1$ corresponds to the robot policy.
III-A Overview
@@
Our approach is based on two insights. The first is that the trajectory of the hand pose in the human demonstration $\mathbf{D}$ provides useful information of the approximate trajectory the robot end-effector should follow, and we can perform kinematic retargeting of the hand trajectory $\{\mathbf{h}_{t}\}^{T}_{t=0}$ to an open-loop robot end-effector trajectory $\{\hat{\mathbf{a}}_{t}\}^{T}_{t=0}$. The second insight is that the kinematically retargeted robot trajectory has the correct form of motion, but these actions may not be very precise in the distribution of plausible robot actions given the current observation. A diffusion policy models this likelihood, and we can use a pre-trained ‘generalist’ diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ to refine the retargeted robot actions in a closed loop manner, thus inferring actions that, while still similar to the human demonstration, are more likely under the policy and better aligned with the robot embodiment (see Fig. 2). We describe both these steps in detail below.
@@
III-B Kinematic retargeting of Human Hand Trajectories to Robot End-Effector Poses
@@
III-C Closed-Loop Denoising of Robot Actions with a Pre-Trained Diffusion Policy
A diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ [10, 45] pre-trained on diverse offline robot interaction data reliably models the distribution of plausible robot actions $\mathbf{a}_{t}$ given previous observations $\mathbf{o}_{\leq t}$. To predict actions $\mathbf{a}_{t}$, diffusion policies start with Gaussian noise $\tilde{\mathbf{a}}_{t}^{(S)}\sim\mathbf{N}(0,I)$ and iteratively denoise it at different diffusion steps $s$:
@@ III-C Closed-Loop Denoising of Robot Actions with a Pre-Trained Diffusion Policy
$\tilde{\mathbf{a}}_{t}^{(s^{*})}=\sqrt{\alpha_{s^{*}}}\hat{\mathbf{a}}_{t}+\sqrt{1-\alpha_{s^{*}}}\epsilon_{t},\quad\epsilon_{t}\sim\mathbf{N}(0,I)$ (2)
Here $\alpha$ corresponds to the diffusion schedule of the pre-trained policy. This procedure, inspired by SDEdit [37] which adopted a similar approach for image editing, relies on the assumption that $\{\hat{\mathbf{a}}_{t}\}_{t=0}^{T}$ is an approximate version of the ideal action trajectory that should be executed by the robot, and it potentially lies outside the distribution of feasible actions under the pre-trained diffusion policy $\bar{\pi}_{\theta}$. Intuitively, adding moderate noise to the retargeted trajectory and then denoising preserves the high-level structure of the human demonstration while allowing the policy to correct low-level infeasibilities from the embodiment gap. The diffusion policy $\bar{\pi}_{\theta}$ then performs iterative denoising steps, conditioned on the robot’s observations $o_{\leq t}$, to refine this noisy trajectory into feasible robot actions, based on equation (2).
After $s^{*}$ denoising steps, the final output $\mathbf{a}_{t}=\tilde{\mathbf{a}}_{t}^{(0)}$ is deployed on the robot. Importantly, this process is carried out in a closed-loop manner: the policy uses real-time observations from the cameras in the scene to iteratively improve its predictions, thereby compensating for embodiment mismatch and external perturbations (e.g., object slippage or occlusion). The key hyperparameter in this process for DemoDiffusion is the diffusion step $s^{*}$ that trades off between the faithfulness to the demonstration and the likelihood under the robot policy – in the limit $s^{*}=S$ we recover a rollout from the base policy $\bar{\pi}_{\theta}$ and in the limit $s^{*}=0$ we recover the kinematically retargeted trajectory $\{\hat{\mathbf{a}}_{t}\}_{t=0}^{T}$. While derived in the context of diffusion, note that this iterative refinement perspective also applies to flow-matching [30], which we show in real-world experiments.
Group Robot Policy Kinematic DemoDiffusion
Small 25.4 2.6 31.8
Medium 26.6 0.7 31.6
Large 27.4 1.5 29.6
Average 26.5 1.6 31.0
TABLE I: Simulated Dexterous Grasping Results. The numbers show the average of success rates over 3 seeds. We use $s^{*}/S=0.2$, where the robot policy uses S=1000S=1000DemoDiffusion consistently outperforms both baselines, across all groups.
Figure 4: Real-World Manipulation Tasks. Human demonstrations for the 8 evaluation tasks, shown as two frames per task. Tasks span prehensile and non-prehensile manipulation, including grasping, pushing, closing, wiping, and placing.
IV Experiments
We evaluate our approach on dexterous grasping in simulation and across diverse real-world tasks comprising prehensile and non-prehensile table-top manipulation. Through experiments, we aim to understand the following research questions: